# Confidence-Based Annotation Of Brain Tumours In Ultrasound

### v1

@@ Abstract
Purpose: An investigation of the challenge of annotating discrete segmentations of brain tumours in ultrasound, with a focus on the issue of aleatoric uncertainty along the tumour margin, particularly for diffuse tumours. A segmentation protocol and method is proposed that incorporates this margin-related uncertainty while minimising the interobserver variance through reduced subjectivity, thereby diminishing annotator epistemic uncertainty. Approach: A sparse confidence method for annotation is proposed, based on a protocol designed using computer vision and radiology theory. Results: Output annotations using the proposed method are compared with the corresponding professional discrete annotation variance between the observers. A linear relationship was measured within the tumour margin region, with a Pearson correlation of 0.8. The downstream application was explored, comparing training using confidence annotations as soft labels with using the best discrete annotations as hard labels. In all evaluation folds, the Brier score was superior for the soft-label trained network. Conclusion: A formal framework was constructed to demonstrate the infeasibility of discrete annotation of brain tumours in B-mode ultrasound. Subsequently, a method for sparse confidence-based annotation is proposed and evaluated. Keywords: Brain tumours, ultrasound, confidence, annotation.
1Hamlyn Centre, Imperial College London, UK
2Department of Imaging, Charing Cross Hospital, UK
3Department of Neurosurgery, Charing Cross Hospital, UK
@@ 1 Introduction
Figure 1: Tumour margin delineation uncertainty among four professional annotators. The three vectors show visible ambiguity over boundary point.
The annotation of objects of interest in medical images is invaluable from both a clinical and technical perspective. For normal anatomy or pathological tissue, the annotation of these features can serve a multitude of purposes that can include helping to make treatment decisions (for example, long-term monitoring) [1], or postoperative evaluation (for example, the extent of resection)[2]. In clinical practice, tumours can be annotated in many ways, such as diameter measurement [3] or arrows for localisation [4]. The most precise form of annotation for differentiating pathological tissue from normal tissue is discrete segmentation [5]. This involves the binary separation of an object from its surroundings/background along the boundary of the object or separation via different categories of the tumour and surrounding normal tissue, such as enhancing, non-enhancing, oedema and necrosis.
In the clinical setting, ultrasound (US) is one of the most widely used imaging modalities. The NHS publication “Diagnostic Imaging Dataset Annual Statistical Release 2022/23” [6] reports that between the years 2012/13 to 2022/23, US imaging accounted for $>20\%$ of all recorded imaging activity, the second most commonly used modality behind X-rays, where the common application of US was for the abdomen/pelvis.
@@
Intraoperative ultrasound (IOUS) has become a commonly used tool to assist in surgical resection, by localising pathological tissue and delineating the margins [10]. Recent evidence suggests that IOUS might be particularly useful in improving the likelihood of maximal safe resection during brain tumour surgery [11]. Although a useful tool, due to the limited standardisation of training and the difficulty of interpreting images due to the confined field of view, the annotation of brain tumours in IOUS images can be difficult and result in variance between observers [12]. This is further compounded in the case of diffuse tumours, where there is often an extremely poor separability of the tumour from the normal surrounding tissue due to the graduated infiltrative nature of the disease and the inherent lack of a boundary [13, 14, 15, 16, 17].
This article explores probabilistic annotation of brain tumours as an alternative to discrete annotation. Ill-definition, diffusion, and infiltration of a tumour and its margin ultimately result in an incompatibility with discrete annotation. What is argued in this article is that annotation should be performed not on a pixel level but rather spatially and locally. The key contributions are as follows.
  • The demonstration of the issue of discrete segmentation of brain tumours in US, where a logic framework is constructed to present this concern formally.
  • A confidence-based annotation protocol and method is proposed.
  • Training tumour segmentation networks with clinically relevant soft labels improves neural network calibration compared to training with hard labels.
@@ 2 Methodology
2.1 The Problem Of Discrete Annotation Of Tumours In US
First, the issue of discrete annotation is addressed. This subsection briefly addresses the logical incoherence of the standard method for annotating diffuse brain tumours on IOUS images. A simple and general logical framework is presented to formalise the problem addressed in this article.
A binary segmentation tasks $S$ will contain an object of interest $O$ within an image $I$. The output of the segmentation task $S(O,I)$ creates an output binary mask $M(O,I)$. Where each pixel $x,y\in M(O,I)$ has the value of $1$ or 00 (True or False). For this to be achievable, the object to be segmented needs to inherently have defined boundaries.
$C(O,I)\Rightarrow S(O,I)$ (1)
Here, $C(O,I)$ defines the condition that the object can be discretely separated from the rest of the image. If the object does not meet this condition (i.e., highly infiltrative diffuse brain tumours $B(O)$ in US image $U(I)$), then it can be expected that the annotation of the object will be affected by aleatoric uncertainty $\Delta_{A}$
$\neg C(O,I)\Rightarrow S(O,I)=\mathbb{E}[\Delta_{A}(M(O,I))]$ (2)
Where the aleatoric uncertainty can be viewed as producing a probability distribution $P$ over the boundary of the object
$\Delta_{A}(M(O,I))=P((x,y)\in\textit{boundary}\ |O,I)$ (3)
Therefore, we use this to define that for brain tumours in US, particularly diffuse tumours, discrete segmentation cannot be performed.
$\neg(B(O)\wedge U(I)\Rightarrow S(O,I))$ (4)
As such, we propose a method for annotation that aims to minimise the impact of aleatoric uncertainty attributed to the fuzzy tumour margins. This approach is designed to reduce subjectivity, thus decreasing interobserver variance and epistemic uncertainty at the annotator level.
@@ 2.2 Annotation Protocol
For the construction of the confidence map of the object, the annotator can use three criteria of scatter points - high, medium, and low confidence. Each scatter point is used to delimit the edge of the corresponding level of confidence over which a heat map is to be constructed. Based on computer vision and radiological theory, a protocol for defining these points has been constructed. Inspiration has been taken from radiological annotation approaches that differentiate the different layers of the tumour and the relevant surrounding tissue, as was done in [5]. Principally, this protocol encourages sensitivity over specificity due to possible extreme levels of aleatoric uncertainty along the tumour margins.
@@
Areas of high confidence can be perceived as where the tumour mass is well defined with expected low levels of interobserver annotation variance - low aleatoric uncertainty in the pixel-level features. Where the granularity of US is sufficient to make the decision on whether to resect. This can be contributed to the area belonging to the hypercellular components of the tumour, which is usually hyperechogenic and thus considerably brighter than it’s surroundings. Furthermore, where the texture is anomalous, significantly homogeneous or heterogeneous relative to the surrounding normal tissue. Or, where edges are well defined by calcification or necrosis.
@@
The medium confidence annotation should primarily be associated with the edge of the fuzzy margin. Here there is likely a large degree of interobserver variability, and the margin aleatoric uncertainty is already high. For these zones, the decision to resect is difficult to make using purely the US image as a reference, where granularity begins to become or is unreliable. In particular, for a diffuse tumour, this is somewhere along the gradient between a clear tumour core and perceived normal tissue. For a well-defined tumour, this can be areas around the clear margin of the tumour that have the continued visual properties of the tumour. For example, continued brightness and or continued anomalous texture. This can also be due to obscuration or acoustic shadowing. This level of confidence is likely the hardest to annotate.
@@
The distribution of low-confidence annotations will occur sparsely outside of the clear tumour mass and the “clear fuzzy margin”. The margin aleatoric uncertainty and interobserver annotation variance are significant. The US granularity is not sufficient to make actionable decisions on resection. These areas will likely reflect the surrounding abnormal tissue that is a variable combination of oedema and less cellular tumour infiltration that correlates with the perilesional, non-enhancing abnormal signal observed on MRI [18]. Pragmatically, these would be the areas that would not typically be resected, as it is thought to reflect predominantly oedema, but where there remains an uncertainty and risk of residual tumour cells.
@@ 2.3 Confidence Map Construction
Once the annotation is complete, the following sets of coordinates are created: high likelihood $\mathcal{C_{H}}=\{(x_{a},y_{a})\}_{a=1}^{n_{H}}$, medium likelihood $\mathcal{C_{M}}=\{(x_{b},y_{b})\}_{b=1}^{n_{M}}$, and low likelihood $\mathcal{C_{L}}=\{(x_{c},y_{c})\}_{c=1}^{n_{L}}$.
@@
The mesh grid is then weighted by aligning annotation coordinates with grid points.
@@
Where Delaunay triangulation $\mathcal{DT}$ is used to determine which grid points correspond to each likelihood region. $\beta_{H},\beta_{M},\beta_{L}$ was assigned values $1,0.5,0.1$.
The confidence map $\mathcal{CM}$ is finally constructed by applying a Gaussian Kernel Density Estimation (KDE), with normalisation, over the mesh grid $\mathcal{G}$ with a minimal bandwidth factor $\sigma=0.1$ chosen to prevent the over-influence of high confidence and further reduction of spatial specificity.
@@
where $n$ is the number of points in $\mathcal{G}$. Overall, the set of tunable parameters is $[\lambda,\beta_{H},\beta_{M},\beta_{L},\sigma]$, which corresponds to [meshgrid size, annotation class weighting, KDE bandwidth].
@@ 3 Experiments And Results
This section will describe the evaluation setup that has been used to validate the usability of this method. Through, the implicit prediction of interoberserver variance as well as down stream data learning application. Sample annotations are displayed in Figure. 2.
Figure 2: Sample ground truth annotations.
@@ 3.1.1 Metric Similarity
Soft $\text{Soft Dice}=1-\frac{2\sum_{i=1}^{n}p_{i}y_{i}}{\sum_{i=1}^{n}p_{i}^{2}+\sum_{i=1}^{n}y_{i}^{2}}$ (9)
Where n is the dataset image index. The Brier score is calculated for both the entire image (considering foreground and background) and exclusively for the foreground, where the foreground is defined as the union of the agreement and confidence map. The Brier score with background is 0.00762 and without 0.0437. The Soft Dice is calculated on the entire image, and an average score of 0.076 is recorded. Both metrics indicate a strong agreement between the confidence map and Agreement maps.
@@ 3.1.2 Similarity Across Threshold Levels
Figure 3: Mean Receiver Operator Characteristic (ROC) curves across threshold interobserver agreement levels. Left subfigure shows the ROC will all pixels included, while the right subfigure shows the ROC only using union area between the two annotations.
To further assess how well the confidence maps align with the Agreement maps. The Receiver Operating Characteristic (ROC) curve are computed at four different binary thresholds of the ground truth, $[0.25,0.50,0.75,1.00]$. The ROC for across the dataset is calculated for both the entire image and exclusively for the foreground. The results of this ROC analysis are shown in Figure. 3. For both cases, the Area Under the Curve (AUC) is greater than $0.9$, with the AUC improving as the threshold increases, which can be attributed to the reduced margin uncertainty. This result adds further evidence that the proposed protocol for creating confidence maps is effective in estimating interobserver variance.
@@
$\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$ (10)
$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (11)
[…]
$\text{Accuracy}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}$ (13)
Where TP is True Positives, TN is True Negatives, FP is False Positives, and FN is False Negatives. The results of this experiment are shown in Figure. 4, where is is notable that all metrics highlight good similarity except at the $1.0$ threshold. Where there is reduced performance in Recall and F1-score. This can be attributed with an increased number of false negatives, due to the KDE’s smoothness and therefore a reduced number of ones.
@@ 3.2 Margin Uncertainty Correlation
Extending the previous evaluation based on binarisation and segment analysis. Analysis of the percentage of non-1 values in both confidence and interoberserver variance is conducted, focused on the margin uncertainty correlation between confidence map and Agreement map.
$\psi_{\text{conf}}$ =|$=\frac{\left|\{(x,y)\in I\mid 0<\text{conf}(x,y)<1\}\right|}{\left|\{(x,y)\in I\mid 0<\text{conf}(x,y)\}\right|}\times 100$
$\psi_{\text{agree}}$ =|$=\frac{\left|\{(x,y)\in I\mid 0<A(x,y)<1\}\right|}{\left|\{(x,y)\in I\mid 0<A(x,y)\}\right|}\times 100$
The metrics for this analysis are the Pearson correlation coefficient and Spearman’s rank correlation coefficient.
@@
The results are displayed in Figure. 5. Where the measured Pearson correlation was $0.80$ and the Spearman correlation $0.75$. The scatter plot shows the correlation between confidence map and interobserver variance, with a linear regression model fitted to the data with a 95% confidence interval. The conclusion of this experiment is that the proposed annotation method will correctly increase the spread of confidence as the size of the tumour and its margins (and uncertainty) increase spatially.
What can be concluded from this set of evaluations is that there is a correlation between confidence maps and unseen interobserver variance. Not only does this affirm the benefit of this protocol and method of annotation in minimising the annotator-level epistemic uncertainty through treating the annotation as a local density spatial task rather than at a discrete pixel/voxel level. But also, these results suggest that this annotation strategy would greatly reduce time taken, as there is a reduce or even no requirement for multiple annotators.
@@ 3.3 Downstream Application: Soft Label Learning
The proposed annotation method is faster and simpler than traditional segmentation and can facilitate dataset generation. Confidence annotations, which cover a larger pixel area and incorporate likelihood grading, may provide more nuanced information about tumours. In deep learning segmentation tasks, probabilistic pixel-wise classification is typically performed using sigmoid or softmax activation functions, with discrete annotations.
Studies, such as [19, 20, 21] have explored the benefit of using soft labels, continuous labels in the binary case, and non-one-hot vectors in the multiclass, for regularisation. Where [22] found that soft labels improved the sensitivity of segmentation networks, particularly for small objects in medical images.
Experiments are performed by training a standard UNet model on both the confidence (soft) and discrete (hard) annotations, where the latter was the annotation from the neuroradiologist. Individually trained to assess the downstream benefit of this method for annotation.
For this experiment, a UNet model with 4 encoder/decoder layers and a maximum channel depth of 64 was trained using sigmoid activation in the final layer. Training was performed separately for each using the Brier and Soft Dice (Equation. 89), as well as the following loss functions, under a 5-fold cross-validation framework:
Binary Cross $\text{Binary Cross Entropy}=-\frac{1}{n}\sum_{i=1}^{n}\left[y_{i}\log(p_{i})+(1-y_{i})\log(1-p_{i})\right]$ (16)
@@
[…]

### v2

@@ Abstract
Purpose: An investigation of the challenge of annotating discrete segmentations of brain tumours in ultrasound, with a focus on the issue of aleatoric uncertainty along the tumour margin, particularly for diffuse tumours. A segmentation protocol and method is proposed that incorporates this margin-related uncertainty while minimising the interobserver variance through reduced subjectivity, thereby diminishing annotator epistemic uncertainty. Approach: A sparse confidence method for ultrasound data annotation is proposed, based on a protocol designed using computer vision and radiology theory. Results: Output annotations using the proposed method are compared with the corresponding discrete annotation variance between the annotators. A linear relationship was measured within the tumour margin region, with a Pearson correlation of 0.8. The downstream application of image segmentation was explored, comparing training using confidence annotations as soft labels with using the best discrete annotations as hard labels. In all evaluation folds, the Brier score was superior for the soft-label trained network. Conclusion: A formal framework was constructed to demonstrate the infeasibility of discrete annotation of brain tumours in B-mode ultrasound images. Subsequently, a method for sparse confidence-based annotation is proposed and evaluated.
keywords:
brain tumours, ultrasound, confidence, annotation
[…]
@@ 1 Introduction
The annotation of objects of interest in medical images is invaluable from both a clinical and technical perspective. For normal anatomy or pathological tissue, the annotation of these areas can serve a multitude of purposes that can include helping to make treatment decisions (for example, long-term monitoring) [1], or postoperative evaluation (for example, the extent of resection)[2]. In clinical practice, tumours can be annotated in many ways, such as diameter measurement [3] or arrows for localisation [4]. The most precise form of annotation for differentiating pathological tissue from normal tissue is discrete segmentation [5]. This involves the binary separation of the tumour from its surroundings/background along the boundary of the tumorous area or separation via different categories of the tumour and surrounding normal tissue, such as enhancing, non-enhancing, oedema and necrosis.
In the clinical setting, ultrasound (US) is one of the most widely used imaging modalities. The NHS publication “Diagnostic Imaging Dataset Annual Statistical Release 2022/23” [6] reports that between the years 2012/13 to 2022/23, US imaging accounted for $20\%$ of all recorded imaging activity, the second most commonly used modality behind X-rays, where the common application of US was for the abdomen/pelvis.
Figure 1: Tumour margin delineation variability among four professional annotators. The three vectors (red, green, blue) show visible ambiguity over the decision of the edge of the tumour, where the segments are shown with yellow dotted lines of varying transparency.
@@
Intraoperative ultrasound (IOUS) has become a commonly used tool to assist in surgical resection, by localising pathological tissue and delineating its margins [10]. Recent evidence suggests that IOUS might be particularly useful in improving the likelihood of maximal safe resection during brain tumour surgery [11]. Although a useful tool, due to the limited standardisation of specific clinical training and the difficulty of interpreting images due to the confined field of view, the annotation of brain tumours in IOUS images can be difficult, resulting in variance between observers [12]. This is further compounded in the case of diffuse tumours, where there is often an extremely poor separability of the tumour from the normal surrounding tissue due to the graduated infiltrative nature of the disease and the inherent lack of a boundary [13, 14, 15, 16, 17].
This article explores probabilistic annotation of brain tumours as an alternative to discrete annotation. Diffusion, and infiltration of a tumour and ill-definition of its margin ultimately make discrete annotation impossible. What is argued in this article is that annotation should be performed not on a pixel level but rather spatially and locally, in the form of distributed continuous confidence values. The key contributions of this work are as follows:
  • We propose a confidence-based annotation protocol and method.
  • We verify that training tumour segmentation networks with clinically relevant soft labels improves neural network calibration compared to training with hard labels.
@@ 2.1 Annotation Protocol
Confidence maps are the proposed alternative annotation method, compared to standard segmentation. Using heat map distributions to represent the perceptual likelihood/confidence in the tumour. This section describes a guideline for how annotators should interpret images and perform the annotation. Fundamental inspiration has been taken from other radiological annotation approaches that differentiate the different layers of the tumour and the relevant surrounding tissue [5]. Specifically, for the construction of the confidence map of the object, the annotation utilises three grades of scatter points - high, medium, and low confidence. Each scatter point is used to delimit the edge of the corresponding level of confidence over which a heat map is to be constructed/fitted. This protocol encourages sensitivity over specificity of the annotation to account for the the impact of aleatoric uncertainty along the tumour margins. Fig. 2 displays examples of ground truth annotation using the graded scatter points.
@@
Areas of high confidence can be perceived as where the tumour mass is well defined with expected low levels of interobserver annotation variance - low aleatoric uncertainty in the pixel-level. In these areas, the granularity of US is sufficient to make the decision on whether to resect or not. This can be contributed to the area belonging to the hypercellular components of the tumour, which is usually hyperechogenic and thus considerably brighter than it’s surroundings. Furthermore, where the texture is anomalous, significantly homogeneous or heterogeneous relative to the surrounding normal tissue. Or, where edges are well defined by calcification or necrosis.
@@
The medium confidence annotation should primarily be associated with the edge of the fuzzy margin. Here there is likely a large degree of interobserver variability, and the aleatoric uncertainty is already high. For these zones, the decision to resect is difficult to make using purely the US image as a reference, where granularity begins to become or is unreliable. In particular, for a diffuse tumour, this is somewhere along the gradient between a clear tumour core and perceived normal tissue. For a well-defined tumour, this can be areas around the clear margin of the tumour that have the continued visual properties of the tumour. For example, continued brightness and or continued anomalous texture. This can also be due to obscuration or acoustic shadowing. This level of confidence is likely the hardest to annotate.
@@
The distribution of low-confidence annotations will occur sparsely outside of the clear tumour mass and the “clear fuzzy margin”. The aleatoric uncertainty and interobserver annotation variance are significant in these areas. The US granularity is not sufficient to make actionable decisions on resection. These areas will likely reflect the surrounding abnormal tissue that is a variable combination of oedema and less cellular tumour infiltration that correlates with the perilesional, non-enhancing abnormal signal observed on MRI [18]. Pragmatically, these would be the areas that would not typically be resected, as it is thought to reflect predominantly oedema, but where there remains an uncertainty and risk of residual tumour cells.
@@ 2.2 Confidence Map Construction
Once the annotation is complete, the following sets of annotation points are defined: high likelihood $\mathcal{C_{H}}=\{(x_{a},y_{a})\}_{a=1}^{n_{H}}$, medium likelihood $\mathcal{C_{M}}=\{(x_{b},y_{b})\}_{b=1}^{n_{M}}$, and low likelihood $\mathcal{C_{L}}=\{(x_{c},y_{c})\}_{c=1}^{n_{L}}$.
@@
The mesh grid is then weighted by aligning annotation points with grid points.
@@
where, Delaunay triangulation $\mathcal{DT}$ is used to determine which grid points correspond to each likelihood region. $\beta_{H},\beta_{M},\beta_{L}$ were assigned values $1,0.5,0.1$, respectively. These parameters are tunable, the chosen values reflecting the likelihood weighting boundary range (0–1).
The confidence map $\mathcal{CM}$ is finally constructed by applying Gaussian Kernel Density Estimation (KDE) (defined by $f$), followed by min-max normalisation, over the mesh grid $\mathcal{G}$. A minimal bandwidth factor $\sigma=0.1$ was chosen to prevent the over-influence of high-confidence areas and further reduction of spatial specificity.
@@
where, $n$ is the number of points in $\mathcal{G}$. Overall, the set of tunable parameters is $\lambda,\beta_{H},\beta_{M},\beta_{L},\sigma]$, which corresponds to [meshgrid size, annotation class weighting, KDE bandwidth].
@@ 3 Experiments And Results
This section will describe the evaluation setup that has been used to validate the usability of this method. Through, the implicit prediction of interobserver variance as well as down stream data learning application. Sample annotations are displayed in Figure. 2.
Figure 2: Sample ground truth annotations. The green points are points of high confidence, yellow are medium and blue are low.
@@ 3.1.1 Metric Similarity
$\text{Soft Dice}=1-\frac{2\sum_{i=1}^{n}p_{i}y_{i}}{\sum_{i=1}^{n}p_{i}^{2}+\sum_{i=1}^{n}y_{i}^{2}}$ (5)
where, n is the dataset image index. The Brier score is calculated for both the entire image (considering foreground and background) and exclusively for the foreground, where the foreground is defined as the union of the agreement and confidence map. The Brier score with background is 0.00762 and without 0.0437. The Soft Dice is calculated on the entire image, and an average score of 0.076 is recorded. Both metrics indicate a strong agreement between the confidence map and Agreement maps.
@@ 3.1.2 Similarity Across Threshold Levels
Figure 3: Mean Receiver Operator Characteristic (ROC) curves across threshold interobserver agreement levels. Left subfigure shows the ROC with all pixels included, while the right subfigure shows the ROC only using union area between the two annotations.
To further assess how well the confidence maps align with the Agreement maps, the Receiver Operating Characteristic (ROC) curve is computed at four different binary thresholds of the ground truth, $[0.25,0.50,0.75,1.00]$. The ROC for across the dataset is calculated for both the entire image and exclusively for the foreground. The results of this ROC analysis are shown in Figure. 3. For both cases, the Area Under the Curve (AUC) is greater than $0.9$, with the AUC improving as the threshold increases, which can be attributed to the reduced margin uncertainty. This result adds further evidence that the proposed protocol for creating confidence maps is effective in estimating interobserver variance.
@@
The results of this experiment are shown in Figure. 4, where is is notable that all metrics highlight good similarity except at the $1.0$ threshold, where there is reduced performance in Recall and F1-score. This can be attributed to an increased number of false negatives, due to the KDE’s smoothness and therefore a reduced number of ones.
@@ 3.2 Margin Uncertainty Correlation
Extending the previous evaluation based on binarisation and segment analysis, analysis of the percentage of non-1 values in both confidence and interoberserver variance is conducted, focused on the margin uncertainty correlation between confidence map and Agreement map.
$\psi_{\text{conf}}$ $=\frac{\left|\{(x,y)\in I\mid 0<\text{conf}(x,y)<1\}\right|}{\left|\{(x,y)\in I\mid 0<\text{conf}(x,y)\}\right|}\times 100$
$\psi_{\text{agree}}$ $=\frac{\left|\{(x,y)\in I\mid 0<A(x,y)<1\}\right|}{\left|\{(x,y)\in I\mid 0<A(x,y)\}\right|}\times 100$
The metrics for this analysis are the Pearson correlation coefficient ($r$) and Spearman’s rank correlation coefficient ($\rho_{s}$).
@@
The results are displayed in Figure. 5, where the measured Pearson correlation was $0.80$ and the Spearman correlation $0.75$. The scatter plot shows the correlation between confidence map and interobserver variance, with a linear regression model fitted to the data with a 95% confidence interval. The conclusion of this experiment is that the proposed annotation method will correctly increase the spread of confidence as the size of the tumour and its margins (and uncertainty) increase spatially.
What can be concluded from this set of evaluations is that there is a correlation between confidence maps and unseen interobserver variance. Not only does this affirm the benefit of this protocol and method of annotation in minimising the annotator-level epistemic uncertainty through treating the annotation as a local density spatial task rather than at a discrete pixel/voxel level. But also, these results suggest that this annotation strategy would greatly reduce time taken, as there is a reduced or even no requirement for multiple annotators.
@@ 3.3 Downstream Application: Soft Label Learning
The proposed annotation method is faster and simpler than traditional segmentation and can facilitate dataset generation. Confidence annotations, which cover a larger pixel area and incorporate likelihood grading, may provide more nuanced information about tumours. In deep learning segmentation tasks, probabilistic pixel-wise classification is typically performed using sigmoid or softmax activation functions, with discrete annotations. Studies, such as [19, 20, 21] have explored the benefit of using soft labels, continuous labels in the binary case, and non-one-hot vectors in the multiclass case, for regularisation. The work proposed in [22] found that soft labels improved the sensitivity of segmentation networks, particularly for small objects in medical images.
In this work, experiments are performed by training a standard UNet model on both the confidence (soft) and discrete (hard) annotations, where the latter was the annotation from the neuroradiologist. This was done to determine the downstream benefit of this method for annotation against binary segmentation for the application of standard neural network training. For this experiment, a UNet model with 4 encoder/decoder layers and a maximum channel depth of 64 was trained using sigmoid activation in the final layer. Training was performed separately with the Brier and Soft Dice (Equation. 45) loss functions, as well as the following other loss functions, under a 5-fold cross-validation framework:
$\text{Binary Cross Entropy}=-\frac{1}{n}\sum_{i=1}^{n}\left[y_{i}\log(p_{i})+(1-y_{i})\log(1-p_{i})\right]$ (8)
@@
where, the Tversky parameters are $\alpha=0.3,\beta=0.7$ [23]. The metrics used for comparison are the same as in Section. 3.1.1, where the Brier score is particularly informative of calibration performance [24]. Results of this are shown in Table. 1. Displayed in Figure. 6 is a sample of outputs for the Binary Cross Entropy test set. Qualitatively, the soft trained network produces smoother confidence. Notably, more likely to activate with low confidence in high uncertainty regions. Unlike the hard trained network which tends to behave more discretely. Notably, the necrotic sphere in the bottom row is ignored by the hard trained network but low confidence is activated with the soft trained network. Which is a more desirable when considering that the necrotic sphere, in both networks, is surrounded by activated pixels.